OpenStudy (anonymous):

sole the equation for x 10+3x=-26 x= ? can someone explain this to me plz

OpenStudy (anonymous):

10+3x=-26 10-10+3x=-26-10 3x=-36..now divide by 3 on both the sides

OpenStudy (compassionate):

10+3x=-26 3x = -36 x = -36/3...